OpenStudy (anonymous):

y varies inversely as three times x. When x = 6, y = 4. Find x when y = 8 okay so I'm not quite sure, but would it end up being x=3?

OpenStudy (anonymous):

x=12

OpenStudy (lgbasallote):

actually...x = 3 is right :)

OpenStudy (anonymous):

Not even one of the choices. & lol thanks!

OpenStudy (lgbasallote):

\[y = \frac{k}{3x}\] \[4 = \frac{k}{3(6)}\] \[4 = \frac{k}{18}\] \[k = 72\] when y = 8 \[y = \frac{k}{3x}\] \[8 = \frac{72}{3x}\] \[8 = \frac{24}{x}\] \[x = \frac{24}{8}\] \[x = 3\] is this what you did @sambruce0825 ?